hi i need help with this question please and thank you.



X represents the Grade Point Average (GPA) of students at a private college which follows normal distribution with mean of 2.8 and standard deviation of 0.52. Find the percentage of students with GPA between 2.00 to 3.00, P (2< X < 3)


1
Expert's answer
2021-09-01T23:46:48-0400

"\mu=2.8 \\
\sigma= 0.52 \\
P(2<X<3) = P(X<3) -P(X<2) \\
= P(Z< \frac{3-2.8}{0.52}) -P(Z< \frac{2-2.8}{0.52}) \\
= P(Z< 0.384) -P(Z< -1.538)"

Form table of Z Score

"= 0.6495 -0.0620 \\

= 0.5875 \\
= 58.75 \; \%"